OpenStudy (anonymous):

Can someone explain to me how I could solve this: (8 sqrt (6)-2)+(5 sqrt (6)-6 sqrt (2))?

OpenStudy (anonymous):

u mean this \[\sqrt{6}-6\sqrt{2}?\]

OpenStudy (anonymous):

No, it looks like this: \[(8\sqrt{6}-\sqrt{2})+(5\sqrt{6}-6\sqrt{2})\]

OpenStudy (anonymous):

I got \[13\sqrt{6}-6\sqrt{2}\]

OpenStudy (anonymous):

You're off by one square root of 2

OpenStudy (anonymous):

So it would be just a 6 at the end? Not a \[6\sqrt{2}\]

OpenStudy (anonymous):

You made a mistake in\[-\sqrt{2}-6\sqrt{2}\]

OpenStudy (anonymous):

I'm confused...if it's a negative radical, does that mean when it is subtracted, or the next thing is subtracted from it, would that cancel out the radical 2? or would it become radical 4?

OpenStudy (anonymous):

There is a -1 in front of\[\sqrt{2}\]So the question is really -1-6

OpenStudy (anonymous):

And -1-6 is -7?

OpenStudy (anonymous):

You're pretty bright.

OpenStudy (anonymous):

Thanks. :) So then the new answer would then be: \[13\sqrt{6}-7\sqrt{2}\]

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

You have to be confident in your answer. Don't question yourself. Go thee forth, my son.

OpenStudy (anonymous):

Thank you!